OpenStudy (anonymous):

Which of the following equations describes the line shown below? Check all that apply. (4,-6) (-5,3) A. y-5=-(x+3) B. y.4=-(x+6) C. y=-x-1 D. y-3=-(x+5) E.y-6=-(x+4) F. y=-x-2

OpenStudy (anonymous):

I wanna say F

OpenStudy (muzzack):

y2-y1/x2-x1 = m y-y1=m(x-x1)

OpenStudy (anonymous):

hey klip is that just the only answer becaus it said check that all apply

OpenStudy (anonymous):

muzzack what would the answer be

OpenStudy (anonymous):

@Muzzack what would the answer be